OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 5x + 4y = 1 4x + 3y = -1 Answer 9 -7 -9 7 What is the value of the y variable in the solution to the following system of equations? 5x + 4y = 1 4x + 3y = -1 Answer 9 -7 -9 7 @Mathematics

jhonyy9 (jhonyy9):

x+y=2 x=2-y 4(2-y)+3y= -1 8-4y+3y=-1 -y=-9 y=9 x=2-9=-7 x=-7 y=9